OpenStudy (anonymous):

amirs checking account pays 0.75% intrest monthly and has $7.50 service fee.if his balance last month was $825 find the amount of debit or credit to his account after intrest and the service fee are both applied

OpenStudy (anonymous):

solution: (1) 825(0.0075)=6.1875 (2) 6.1875+7.50=13.6875 (3) 825+ 13.6875= 838.6875 or 838.7